OpenStudy (anonymous):

I do need help with this one :D 7= 1/sqrtx

OpenStudy (anonymous):



OpenStudy (anonymous):

x = 1/49

OpenStudy (ash2326):

square both sides we get 1/x=49 so x=1/49